# Xinxu Weia    Kanhao Zhaoa    Yong Jiaoa    Lifang Hea    Yu Zhanga11footnotemark: 1

### v2

@@ Appendix I Comparison of Positional Encoding Methods
We investigate four commonly used positional encoding (PE) strategies for graph neural networks: Laplacian Positional Encoding (LPE), Node Degree Positional Encoding, Brain Gradient Positional Encoding [Dong et al., 2024a], and Random Walk Structural Encoding (RWSE). Below, we define each method and evaluate their characteristics in the context of fMRI-based brain graph modeling.
@@
The attention mask $\mathbf{M}_{i}$ is broadcast across the attention heads and used to mask out the scores corresponding to padded nodes by replacing them with $\infty$- ∞:
@@
A related concept appears in Meta-Matching [He et al., 2022], which aims to transfer pre-trained models from large-scale datasets (e.g., UK Biobank) to small-scale neuroimaging datasets. Meta-Matching freezes the entire backbone and learns a lightweight linear regressor on the model outputs to align with new target labels. Each target dataset is treated as an independent task, without leveraging shared structure or task distributions. While both approaches seek efficient cross-dataset transfer, their adaptation mechanisms and learning paradigms are fundamentally different.
@@
Table 9: Key Differences between Meta-Matching [He et al., 2022] and Our Prompt-based Meta-Learning Framework
@@
Appendix N Details of Graph Masked Autoencoder (GMAE) and Graph Contrastive Learning (GCL)
We propose a unified pre-training framework that combines Graph Masked Autoencoder (GMAE) and Graph Contrastive Learning (GCL) to learn robust and generalizable representations for brain graphs. These two components complement each other: GMAE enables fine-grained feature-level recovery through generative reconstruction, while GCL encourages invariance under perturbations by contrasting different views of the same graph.
N.1 Graph Masked Autoencoder (GMAE)
Inspired by recent advances in masked autoencoding [Hou et al., 2022], we apply random masking to both nodes and edges of the input graph. Given a graph with $N$ nodes, adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, and node features $\mathbf{X}\in\mathbb{R}^{N\times F}$, we randomly sample a subset of nodes $\mathcal{V}_{M}\subset\mathcal{V}$ to mask. For each masked node $v_{i}\in\mathcal{V}_{M}$, its input feature is replaced with a learnable mask token $\mathbf{x}_{[M]}\in\mathbb{R}^{D}$. The masked node feature matrix $\tilde{\mathbf{X}}$ is defined as:
@@
The GMAE encoder processes the corrupted graph $(\tilde{\mathbf{X}},\tilde{\mathbf{A}})$ and produces latent embeddings, which are then passed to a lightweight decoder to reconstruct the original node features of the masked nodes. The reconstruction objective is defined as:
@@
N.2 Graph Contrastive Learning (GCL)
In parallel with the generative pathway, we apply contrastive learning to enforce view-invariant representations. Specifically, we generate two augmented views of the same input graph by applying lightweight stochastic perturbations (i.e., random feature dropout and edge dropout). One view is treated as the query, while the other serves as the key. Let $(\mathbf{X}^{(q)},\mathbf{A}^{(q)})$ and $(\mathbf{X}^{(k)},\mathbf{A}^{(k)})$ denote the two views; these are passed through a shared encoder to obtain corresponding representations $z_{q}$ and $z_{k}$. We adopt the NT-Xent contrastive loss [Qiu et al., 2020] to maximize the similarity between matching query-key pairs from the same graph while distinguishing them from others in the batch:
@@
[…]
@@
Table 10 provides a comprehensive comparison between Graph Contrastive Learning (GCL) and Graph Masked Autoencoder (GMAE) as two prominent pre-training paradigms for graph neural networks. GCL emphasizes learning discriminative representations by contrasting positive and negative graph pairs, making it particularly effective for classification and retrieval tasks. In contrast, GMAE focuses on reconstructing masked parts of the graph, encouraging the model to capture fine-grained structural details and local contextual information. While GCL tends to produce compact and abstract embeddings that distinguish samples globally, GMAE yields richer and more structure-aware representations suitable for reconstruction and local reasoning. However, both approaches have limitations: GCL’s effectiveness heavily depends on the design of graph augmentations, potentially neglecting subtle local cues; GMAE, on the other hand, is sensitive to the masking ratio and may underperform in tasks requiring global discrimination. These complementary characteristics suggest that combining both strategies may lead to more robust and generalizable graph representations.
Table 10: Comparison between Graph Contrastive Learning (GCL) and Graph Masked Autoencoder (GMAE) Pre-training.
Graph Contrastive Learning (GCL) Graph Masked Autoencoder (GMAE)
@@
>5 Rarely Used Risk of overfitting support set and high computational cost. Not commonly used unless thoroughly validated.
@@
Model Foundation Input Type Domain Pre-Training Method Tuning Method Tuning Shot Parcellation Disorder Vanilla GCN Graph Spatial - - - Single Single BrainGNN Graph Spatial - - - Single Single Vanilla TF Connectome/FC Spatial - - - Single Single Graph TF Graph Spatial - - - Single Single BrainNetTF Connectome/FC Spatial - - - Single Single BrainNPT Connectome/FC Spatial Connectome/FC Generative Fine-Tuning Full/Few-shot Single Multiple (< 5) BrainMass Connectome/FC Spatial Connectome/FC Generative Fine-Tuning Full/Few-shot Single Multiple (10) BrainLM Time Series Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (< 5) Brain-JEPA Time Series Spatial+Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (< 5) BrainGFM Graph Spatial Graph (Gener. + Contra.) Fine/Prompt-Tuning Full/Few/Zero-shot Multiple Multiple (25)
The Table 15 provides a systematic comparison of various brain foundation models and baseline methods across multiple dimensions, including architectural type, data domain, pre-training strategy, and tuning method. These approaches can be broadly categorized into three groups: conventional models without pre-training (e.g., Vanilla GCN [Kipf and Welling, 2016], BrainGNN [Li et al., 2021], Vanilla Connectome/FC-based Tansformer (TF) [Yun et al., 2019]), ROI- or time-series-based pre-trained models (e.g., BrainNPT [Hu et al., 2024], BrainMass [Yang et al., 2024], BrainLM [Caro et al., 2023]), and our proposed graph-based foundation model, BrainGFM.
@@
Pre-Train ABCD [Casey et al., 2018] 11,878 35,770 Multiple
ADHD 200 [consortium, 2012] 973 1382 ADHD
ABIDE I [Di Martino et al., 2014] 1,112 1,112 ASD
ADNI 3 [Jack Jr et al., 2008] 1,071 1,410 AD, MCI
AOMIC [Snoek et al., 2021] 210 210 Multiple
AURORA [McLean et al., 2020] 284 284 PTSD
CAM_CAN [Shafto et al., 2014] 652 652 -
CATD [Nielson et al., 2023] 127 454 Multiple
GSP [Holmes et al., 2015] 1,569 2,706 -
HCP-Aging [Bookheimer et al., 2019] 724 724 -
EMBARC [Trivedi et al., 2016] 308 308 MDD
LEMON [Babayan et al., 2019] 213 213 MDD
HABS [Dagley et al., 2017] 284 1,371 -
PREVEND_AD [Tremblay-Mercier et al., 2021] 343 2,427 AD
SRPBS_Japan [Yamashita et al., 2019] 1,410 1,410 ASD
NYU_CUD [Kelly et al., 2011] 29 56 CUD
OASIS3 [LaMontagne et al., 2019] 1,172 4,090 Dementia
HBN [Alexander et al., 2017] 2,228 4,039 Multiple
SubMex_RTMS [Angeles-Valdez et al., 2024] 150 150 CUD
Internal Test ADHD 200 [consortium, 2012] 973 1382 ADHD
HBN [Alexander et al., 2017] 2,282 4,039 Multiple
OASIS3 [LaMontagne et al., 2019] 1,172 4,090 Dementia
Semi-External Test ABIDE II [Di Martino et al., 2014] 1,044 1,044 ASD
ADNI 2 [Weiner et al., 2013] 1,171 1,306 AD, MCI
SubMex_CUD [Angeles-Valdez et al., 2022] 135 135 CUD
External Test UCLA_CNP [Poldrack et al., 2016] 261 261 Multiple
REST-META-MDD [Yan et al., 2019] 2,379 2,379 MDD
@@
For all cohorts, resting-state fMRI data were collected with varying protocols and scanner parameters specific to each study site. All available resting-state fMRI data were preprocessed using the well-established fMRIPrep pipeline [Esteban et al., 2019]. The T1-weighted image was corrected for intensity non-uniformity and then stripped skull. Spatial normalization was done through nonlinear registration, with the T1w reference [Avants et al., 2008]. Using FSL, brain features such as cerebrospinal fluid, white matter, and grey matter were segmented from the reference, brain-extracted T1 weighted image [Zhang et al., 2000]. The fieldmap information was used to correct distortion in low-frequency and high-frequency components of fieldmap. Then, a corrected echo-planar imaging reference was obtained from a more accurate co-registration with the anatomical reference. The blood-oxygenation-level-dependent (BOLD) reference was then transformed to the T1-weighted image with a boundary-based registration method, configured with nine degrees of freedom to account for distortion remaining in the BOLD reference [Greve and Fischl, 2009]. Head-motion parameters (rotation and translation parameters of volume-to-reference transform matrices) were estimated with MCFLIRT (FSL). BOLD signals were slice-time corrected and resampled onto the participant’s original space with head-motion correction, susceptibility distortion’s correction, and then resampled into standard space, generating a preprocessed BOLD run in MNI152NLin2009cAsym space. Automatic removal of motion artifacts using independent component analysis (ICA-AROMA) [Pruim et al., 2015] was performed on the preprocessed BOLD time-series on MNI space after removal of non-steady-state volumes and spatial smoothing with an isotropic Gaussian kernel of 6 mm FWHM (full-width half-maximum).
@@
Category Disease/Disorder Dataset(s) Neurodevelopmental Disorders Attention-Deficit/Hyperactivity Disorder (ADHD) ADHD200 Autism Spectrum Disorder (ASD) ABIDE II Language Disorder (LD) HBN Specific Learning Disorder (SLD) HBN Intellectual Disability (ID) HBN Speech Sound Disorder (SSD) HBN Communication Disorder (CD) HBN Neurodegenerative Disorders Alzheimer’s Disease (AD) ADNI 2 Mild Cognitive Impairment (MCI) ADNI 2 Dementia (DM) OASIS3 Mood and Anxiety Disorders Major Depression Disorder (MDD) HBN, REST-META-MDD Anxiety (ANX) HBN Post-Traumatic Stress Disorder (PTSD) HBN Adjustment Disorder (AJD) HBN Mild Depression Disorder (PDD) HBN Bipolar Disorder (BP) UCLA_CNP Obsessive-Compulsive and Impulse Control Disorders Obsessive-Compulsive Disorder (OCD) HBN Oppositional Defiant Disorder (ODD) HBN Intermittent Explosive Disorder (IED) HBN Motor Disorders Tourette Syndrome (TS) HBN Motor Disorder (MD) HBN Provisional Tic Disorder (PTD) HBN Substance Use Disorders Cocaine Use Disorder (CUD) SubMex_CUD Psychotic Disorders Schizophrenia (SCHZ) UCLA_CNP
@@
Appendix U Broader Impact
Appendix V Limitation and Future Works
While our work successfully constructed a large-scale fMRI dataset for pre-training, certain limitations remain. Due to the significant manual effort involved, we were unable to include all datasets from the OpenNeuro platform [Markiewicz et al., 2021], particularly the large number of task-based (non-resting-state) fMRI datasets. In addition, because of financial constraints, we were not able to incorporate fMRI data from the UK Biobank [Bycroft et al., 2018], including both resting-state and task-based scans, as access to this dataset requires paid licensing and ongoing maintenance costs.

### v3

@@ Appendix I Comparison of Positional Encoding Methods
We investigate four commonly used positional encoding (PE) strategies for graph neural networks: Laplacian Positional Encoding (LPE), Node Degree Positional Encoding, Brain Gradient Positional Encoding (Dong et al., 2024a), and Random Walk Structural Encoding (RWSE). Below, we define each method and evaluate their characteristics in the context of fMRI-based brain graph modeling.
@@
The attention mask $\mathbf{M}_{i}$ is broadcast across the attention heads and used to mask out the scores corresponding to padded nodes by replacing them with -$\infty$:
@@
A related concept appears in Meta-Matching (He et al., 2022), which aims to transfer pre-trained models from large-scale datasets (e.g., UK Biobank) to small-scale neuroimaging datasets. Meta-Matching freezes the entire backbone and learns a lightweight linear regressor on the model outputs to align with new target labels. Each target dataset is treated as an independent task, without leveraging shared structure or task distributions. While both approaches seek efficient cross-dataset transfer, their adaptation mechanisms and learning paradigms are fundamentally different.
@@
Table 9: Key Differences between Meta-Matching (He et al., 2022) and Our Prompt-based Meta-Learning Framework
@@
Appendix N Details of Graph Masked Autoencoder and Graph Contrastive Learning
We propose a unified pre-training framework that combines Graph Masked Autoencoder and Graph Contrastive Learning to learn robust and generalizable representations for brain graphs. These two components complement each other: GMAE enables fine-grained feature-level recovery through generative reconstruction, while GCL encourages invariance under perturbations by contrasting different views of the same graph.
N.1 Graph Masked Autoencoder Pre-Training
Inspired by recent advances in masked autoencoding (Hou et al., 2022), we apply random masking to both nodes and edges of the input graph. Given a graph with $N$ nodes, adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, and node features $\mathbf{X}\in\mathbb{R}^{N\times F}$, we randomly sample a subset of nodes $\mathcal{V}_{M}\subset\mathcal{V}$ to mask. For each masked node $v_{i}\in\mathcal{V}_{M}$, its input feature is replaced with a learnable mask token $\mathbf{x}_{[M]}\in\mathbb{R}^{D}$. The masked node feature matrix $\tilde{\mathbf{X}}$ is defined as:
@@
where $\mathbf{M}_{e}\in\{0,1\}^{N\times N}$ is a symmetric binary mask sampled from a Bernoulli distribution with parameter 1p1-p, and $\odot$ denotes element-wise multiplication.
@@
N.2 Graph Contrastive Learning Pre-Training
In parallel with the generative pathway, we apply contrastive learning to enforce view-invariant representations. Specifically, we generate two augmented views of the same input graph by applying lightweight stochastic perturbations (i.e., random feature dropout and edge dropout). One view is treated as the query, while the other serves as the key. Let $(\mathbf{X}^{(q)},\mathbf{A}^{(q)})$ and $(\mathbf{X}^{(k)},\mathbf{A}^{(k)})$ denote the two views; these are passed through a shared encoder to obtain corresponding representations $z_{q}$ and $z_{k}$. We adopt the NT-Xent contrastive loss (Qiu et al., 2020) to maximize the similarity between matching query-key pairs from the same graph while distinguishing them from others in the batch:
@@
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature hyperparameter, and $B$ is the batch size. Here, $b$ and $b^{\prime}$ index different samples within the batch, where each sample corresponds to an augmented graph.
@@
Table 10 provides a comprehensive comparison between Graph Contrastive Learning and Graph Masked Autoencoder as two prominent pre-training paradigms for graph neural networks. GCL emphasizes learning discriminative representations by contrasting positive and negative graph pairs, making it particularly effective for classification and retrieval tasks. In contrast, GMAE focuses on reconstructing masked parts of the graph, encouraging the model to capture fine-grained structural details and local contextual information. While GCL tends to produce compact and abstract embeddings that distinguish samples globally, GMAE yields richer and more structure-aware representations suitable for reconstruction and local reasoning. However, both approaches have limitations: GCL’s effectiveness heavily depends on the design of graph augmentations, potentially neglecting subtle local cues; GMAE, on the other hand, is sensitive to the masking ratio and may underperform in tasks requiring global discrimination. These complementary characteristics suggest that combining both strategies may lead to more robust and generalizable graph representations.
Table 10: Comparison between Graph Contrastive Learning and Graph Masked Autoencoder Pre-training.
Graph Contrastive Learning - GCL Graph Masked Autoencoder - GMAE
@@
5 Rarely Used Risk of overfitting support set and high computational cost. Not commonly used unless thoroughly validated.
@@
Model Foundation Input Type Domain Pre-Training Method Tuning Method Tuning Shot Parcellation Disorder Vanilla GCN Graph Spatial - - - Single Single BrainGNN Graph Spatial - - - Single Single Vanilla TF Connectome/FC Spatial - - - Single Single Graph TF Graph Spatial - - - Single Single BrainNetTF Connectome/FC Spatial - - - Single Single BrainNPT Connectome/FC Spatial Connectome/FC Generative Fine-Tuning Full/Few-shot Single Multiple (<< 5) BrainMass Connectome/FC Spatial Connectome/FC Generative Fine-Tuning Full/Few-shot Single Multiple (10) BrainLM Time Series Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (<< 5) Brain-JEPA Time Series Spatial+Temporal Time Series Generative Fine-Tuning Full-shot Single Multiple (<< 5) BrainGFM Graph Spatial Graph (Gener. + Contra.) Fine/Prompt-Tuning Full/Few/Zero-shot Multiple Multiple (25)
The Table 15 provides a systematic comparison of various brain foundation models and baseline methods across multiple dimensions, including architectural type, data domain, pre-training strategy, and tuning method. These approaches can be broadly categorized into three groups: conventional models without pre-training (e.g., Vanilla GCN (Kipf and Welling, 2016), BrainGNN (Li et al., 2021), Vanilla Connectome/FC-based Tansformer (TF) (Yun et al., 2019)), ROI- or time-series-based pre-trained models (e.g., BrainNPT (Hu et al., 2024), BrainMass (Yang et al., 2024), BrainLM (Caro et al., 2023)), and our proposed graph-based foundation model, BrainGFM.
@@
Pre-Train ABCD (Casey et al., 2018) 11,878 35,770 Multiple
ADHD 200 (consortium, 2012) 973 1382 ADHD
ABIDE I (Di Martino et al., 2014) 1,112 1,112 ASD
ADNI 3 (Jack Jr et al., 2008) 1,071 1,410 AD, MCI
AOMIC (Snoek et al., 2021) 210 210 Multiple
AURORA (McLean et al., 2020) 284 284 PTSD
CAM_CAN (Shafto et al., 2014) 652 652 -
CATD (Nielson et al., 2023) 127 454 Multiple
GSP (Holmes et al., 2015) 1,569 2,706 -
HCP-Aging (Bookheimer et al., 2019) 724 724 -
EMBARC (Trivedi et al., 2016) 308 308 MDD
LEMON (Babayan et al., 2019) 213 213 MDD
HABS (Dagley et al., 2017) 284 1,371 -
PREVEND_AD (Tremblay-Mercier et al., 2021) 343 2,427 AD
SRPBS_Japan (Yamashita et al., 2019) 1,410 1,410 ASD
NYU_CUD (Kelly et al., 2011) 29 56 CUD
OASIS3 (LaMontagne et al., 2019) 1,172 4,090 Dementia
HBN (Alexander et al., 2017) 2,228 4,039 Multiple
SubMex_RTMS (Angeles-Valdez et al., 2024) 150 150 CUD
Internal Test ADHD 200 (consortium, 2012) 973 1382 ADHD
HBN (Alexander et al., 2017) 2,282 4,039 Multiple
OASIS3 (LaMontagne et al., 2019) 1,172 4,090 Dementia
Semi-External Test ABIDE II (Di Martino et al., 2014) 1,044 1,044 ASD
ADNI 2 (Weiner et al., 2013) 1,171 1,306 AD, MCI
SubMex_CUD (Angeles-Valdez et al., 2022) 135 135 CUD
External Test UCLA_CNP (Poldrack et al., 2016) 261 261 Multiple
REST-META-MDD (Yan et al., 2019) 2,379 2,379 MDD
@@
For all cohorts, resting-state fMRI data were collected with varying protocols and scanner parameters specific to each study site. All available resting-state fMRI data were preprocessed using the well-established fMRIPrep pipeline (Esteban et al., 2019). The T1-weighted image was corrected for intensity non-uniformity and then stripped skull. Spatial normalization was done through nonlinear registration, with the T1w reference (Avants et al., 2008). Using FSL, brain features such as cerebrospinal fluid, white matter, and grey matter were segmented from the reference, brain-extracted T1 weighted image (Zhang et al., 2000). The fieldmap information was used to correct distortion in low-frequency and high-frequency components of fieldmap. Then, a corrected echo-planar imaging reference was obtained from a more accurate co-registration with the anatomical reference. The blood-oxygenation-level-dependent (BOLD) reference was then transformed to the T1-weighted image with a boundary-based registration method, configured with nine degrees of freedom to account for distortion remaining in the BOLD reference (Greve and Fischl, 2009). Head-motion parameters (rotation and translation parameters of volume-to-reference transform matrices) were estimated with MCFLIRT (FSL). BOLD signals were slice-time corrected and resampled onto the participant’s original space with head-motion correction, susceptibility distortion’s correction, and then resampled into standard space, generating a preprocessed BOLD run in MNI152NLin2009cAsym space. Automatic removal of motion artifacts using independent component analysis (ICA-AROMA) (Pruim et al., 2015) was performed on the preprocessed BOLD time-series on MNI space after removal of non-steady-state volumes and spatial smoothing with an isotropic Gaussian kernel of 6 mm FWHM (full-width half-maximum).
@@
Category Disease/Disorder Dataset(s) Neurodevelopmental Disorders Attention-Deficit/Hyperactivity Disorder (ADHD) ADHD200 Autism Spectrum Disorder (ASD) ABIDE II Language Disorder (LD) HBN Specific Learning Disorder (SLD) HBN Intellectual Disability (ID) HBN Speech Sound Disorder (SSD) HBN Communication Disorder (CD) HBN Neurodegenerative Disorders Alzheimer’s Disease (AD) ADNI 2 Mild Cognitive Impairment (MCI) ADNI 2 Dementia (DM) OASIS3 Mood and Anxiety Disorders Major Depression Disorder (MDD) HBN, REST-META-MDD Anxiety (ANX) HBN Post-Traumatic Stress Disorder (PTSD) HBN Adjustment Disorder (AJD) HBN Mild Depression Disorder (PDD) HBN Bipolar Disorder (BP) UCLA_CNP Obsessive-Compulsive and Impulse Control Disorders Obsessive-Compulsive Disorder (OCD) HBN Oppositional Defiant Disorder (ODD) HBN Motor Disorders Tourette Syndrome (TS) HBN Motor Disorder (MD) HBN Provisional Tic Disorder (PTD) HBN Substance Use Disorders Cocaine Use Disorder (CUD) SubMex_CUD Psychotic Disorders Schizophrenia (SCHZ) UCLA_CNP
@@
Appendix U Brain Graph Interpretability Analysis
As shown in Figure U, we present ROI-level attention attribution maps derived from the graph transformer layers of BrainGFM to illustrate how the model utilizes functional brain graph structures for prediction. We visualize attribution differences between normal control (NC) and patient groups across two datasets: ABIDE (Autism Spectrum Disorder, ASD) and HBN (Major Depressive Disorder, MDD). Each heatmap represents the average pooled attention attribution values across all subjects within the corresponding group. Clear differences in the attribution patterns can be observed between ASD vs. normal controls in ABIDE and MDD vs. normal controls in HBN. Regions with consistently higher attribution values indicate stronger contributions to classification decisions. Notably, several frontal and temporal cortical regions, frequently reported as clinically relevant in psychiatric neuroimaging studies, exhibit salient involvement. The distinct attention distributions demonstrate that BrainGFM captures meaningful disorder-related functional subnetworks rather than relying on noise or dataset-specific artifacts.
These results support the interpretability of the learned graph representations and provide preliminary evidence that BrainGFM can highlight potential biomarkers associated with neuropsychiatric disorders, helping bridge model decisions with neuroscientific insights.
Figure 10: ROI-level attention attribution maps for ABIDE (ASD) and HBN (MDD), showing distinct attention patterns between normal controls and patients.